Help! I need help understanding how to arrive at the correct solution to each section of this particular problem. I would greatly appreciate it.

Before any debt cancellation, the insolvent KuhnCo holds business equipment, its only asset, with a fair market value of $1 million and related liabilities of $1.25 million. The lender agrees to cancel $400,000 of the liabilities. KuhnCo has no other liabilities.

a. How much gross income does KuhnCo report as a result of the debt cancellation?

b. How would your answer change, if at all, had the lender canceled $200,000 of the debt?
Answer #1

IN THE FIRST CASE ,

ASSET VALUE= 1000000

LIABILITY= 1250000

LENDER AGREES TO CANCEL 400000 ie , LIABILITY FALLS TO $850000

SO GROSS INCOME IS 1000000-850000= $150000

SECOND CASE

ASSET VALUE =1000000

LIABILITY AFTER CANCELLATION = 1250000-200000=1050000

GROSS INCOME/LOSS = 1000000-1050000= (-50000) ie, LOSS OF $50000